OpenStudy (anonymous):

HELP PLEASE!!!!!! I know the answer I just need to know about the exponents...

OpenStudy (anonymous):

\[\sqrt{72x ^{3}y ^{16}}\]

Nnesha (nnesha):

you can convert square root of something to 1/2 exponents right \[\huge\rm \sqrt{x} =x^\frac{ 1 }{ 2 }\] what are the factors of 75 ?

Nnesha (nnesha):

x^3 can be written as x^2 times x \[\huge\rm \sqrt{72 · x^2 · x· y^{16}} = \sqrt{72x} · x^\frac{ 2 }{ 2 } y^\frac{ 16 }{ 2 }\]

Nnesha (nnesha):

let me know if you have question about that^^^^

OpenStudy (anonymous):

why do you separate the x^2 and x

Nnesha (nnesha):

so i can take square root of x^2 we have to take out the variables from the square root or another way of taking square root of x^3 would